# Sharing Key Semantics in Transformer Makes Efficient Image Restoration

### v1

@@ A.3 Efficiency Analysis
$\mathcal{O}(6\times[4HWC^{2}+2kHWC]+HWC)$ (5)
@@
$\mathcal{O}(6\times[4HWC^{2}+2kHWC]+HWC-6\times[4HWC^{2}+2(M)^{2}HWC])$ (6)
$=\mathcal{O}((12M^{2}-K-1)HWC)>>0$
In the last equation, since $M^{2}$ contains all the patches/pixels within a certain window, for example, in a common setting with a window size of M = 7 and a patch size of 16, there are around (7 $\times$ 7) $\times$ (16 $\times$ 16) pixels within the given window. However, the k value in our proposed Key-Semantic dictionary is set to 512 or randomly sampled from [64, 128, 256, 384, 512], which is significantly smaller than (7 $\times$ 7) $\times$ (16 $\times$ 16).
@@ A.4 Training Details
Our method explores 6 IR tasks, and the training settings vary slightly for each task. These differences encompass the architecture of the proposed SemanIR, variations in the choice of the optimizer, and loss functions.
@@
Figure 8: Ablation study on the impact of $k$ for Grayscale JPEG CAR on BSD500 (a-d), Classic5 (e-h), Live1(i-l), and Urban100 (m-p) datasets with $QF$ = {10, 20, 30, 40}.
@@ Appendix D More Ablation Analyses
The Impact of the $k$ in Key-Semantic Dictionary Construction under Various IR Tasks. To explore how the $k$ value of top-k will affect the IR performance of the proposed SemanIR. We conduct exhaustive experiments on JPEG compression artifact reduction for both color and grayscale images under different QF values (i.e., QF = [10, 20, 30, 40]), image denoising for both color and grayscale images under different noise levels (i.e., $\sigma$ = [15, 25, 50]), as well as image SR under different scales (i.e., 2$\times$, 3$\times$, 4$\times$) with the proposed SemanIR. Note that all the experiments for each IR task are conducted under two kinds of top-k settings, i.e., (i) $k$ was randomly sampled from the range [64, 512] during the overall training phase, and (ii) $k$ was held constant at 512 throughout the training phase. For inference, $k$ was configured to the specified value for both settings.
The results of the JPEG CAR in terms of the hyper-parameters $k$ under different training settings during inference for both color image and grayscale image are shown in Fig. 7 and Fig. 8. It is clear that for both the color and grayscale JPEG CAR task, when $k$ is set to 64 during inference, there is a huge performance cat between the random top-k setting and the fixed top-k setting. In addition, the fixed top-k setting performs well or sometimes even a bit better than the random top-k setting only when $k$ is also set to the same number (i.e., 512). With the decrease of $k$ during inference for the fixed top-k setting, the PSNR drops largely marginally for all the datasets under every kind of degraded QF factor.
[…]
Table 11: Full Grayscale image JPEG compression artifact removal results.
Set QF JPEG DnCNN3 [104] DRUNet [103] SemanIR (Ours) GRL-S [50] SwinIR [52] ART [102] CAT [16] SemanIR (Ours)
[…]
Classic5 10 27.82 0.7600 29.40 0.8030 30.16 0.8234 30.26 0.8240 30.20 0.8286 30.27 0.8249 30.27 0.8258 30.26 0.8250 30.36 0.8267
20 30.12 0.8340 31.63 0.8610 32.39 0.8734 32.52 0.8740 32.49 0.8776 32.52 0.8748 - - 32.57 0.8754 32.58 0.8748
30 31.48 0.8670 32.91 0.8860 33.59 0.8949 33.74 0.8955 33.72 0.8985 33.73 0.8961 33.74 0.8964 33.77 0.8964 33.77 0.8958
40 32.43 0.8850 33.77 0.9000 34.41 0.9075 34.55 0.9078 34.53 0.9107 34.52 0.9082 34.55 0.9086 34.58 0.9087 34.57 0.9080
LIVE1 10 27.77 0.7730 29.19 0.8120 29.79 0.8278 29.84 0.8323 29.82 0.8323 29.86 0.8287 29.89 0.8300 29.89 0.8295 29.92 0.8360
20 30.07 0.8510 31.59 0.8800 32.17 0.8899 32.23 0.8949 32.22 0.8930 32.25 0.8909 - - 32.30 0.8913 32.28 0.8950
30 31.41 0.8850 32.98 0.9090 33.59 0.9166 33.65 0.9213 33.65 0.9190 33.69 0.9174 33.71 0.9178 33.73 0.9177 33.69 0.9201
40 32.35 0.9040 33.96 0.9250 34.58 0.9312 34.65 0.9329 34.64 0.9331 34.67 0.9317 34.70 0.9322 34.72 0.9320 34.67 0.9345
Urban100 10 26.33 0.7816 28.54 0.8484 30.31 0.8745 30.81 0.8885 30.70 0.8875 30.55 0.8835 30.87 0.8894 30.81 0.8866 31.15 0.8941
20 28.57 0.8545 31.01 0.9050 32.81 0.9241 33.33 0.9266 33.24 0.9270 33.12 0.9190 - - 33.38 0.9269 33.51 0.9272
30 30.00 0.9013 32.47 0.9312 34.23 0.9414 34.74 0.9446 34.67 0.9430 34.58 0.9417 34.81 0.9442 34.81 0.9449 34.84 0.9462
40 31.06 0.9215 33.49 0.9412 35.20 0.9547 35.69 0.9447 35.62 0.9519 35.50 0.9515 35.73 0.9553 35.73 0.9511 35.75 0.9550
@@
Table 12: Full Classical image SR results. Both lightweight and accurate models are summarized.
Method Scale Params Set5 Set14 BSD100 Urban100 Manga109
[…]
NLSA [64] 2$\times$ 42.63 38.34 0.9618 34.08 0.9231 32.43 0.9027 33.42 0.9394 39.59 0.9789
IPT [10] 2$\times$ 115.48 38.37 - 34.43 - 32.48 - 33.76 - - -
RCAN [111] 3$\times$ 15.63 34.74 0.9299 30.65 0.8482 29.32 0.8111 29.09 0.8702 34.44 0.9499
SAN [19] 3$\times$ 15.90 34.75 0.9300 30.59 0.8476 29.33 0.8112 28.93 0.8671 34.30 0.9494
HAN [68] 3$\times$ 64.35 34.75 0.9299 30.67 0.8483 29.32 0.8110 29.10 0.8705 34.48 0.9500
NLSA [64] 3$\times$ 45.58 34.85 0.9306 30.70 0.8485 29.34 0.8117 29.25 0.8726 34.57 0.9508
IPT [10] 3$\times$ 115.67 34.81 - 30.85 - 29.38 - 29.49 - - -
SwinIR [52] 3$\times$ 11.94 34.97 0.9318 30.93 0.8534 29.46 0.8145 29.75 0.8826 35.12 0.9537
CAT-A [16] 3$\times$ 16.64 35.06 0.9326 31.04 0.8538 29.52 0.8160 30.12 0.8862 35.38 0.9546
ART [102] 3$\times$ 16.58 35.07 0.9325 31.02 0.8541 29.51 0.8159 30.10 0.8871 35.39 0.9548
EDT [49] 3$\times$ 11.66 35.13 0.9328 31.09 0.8553 29.53 0.8165 30.07 0.8863 35.47 0.9550
SemanIR-S (Ours) 3$\times$ 12.05 34.99 0.9366 31.23 0.8594 29.53 0.8223 30.71 0.8950 35.52 0.9573
SemanIR-B (Ours) 3$\times$ 20.08 35.03 0.9371 31.29 0.8603 29.54 0.8227 30.87 0.9012 35.60 0.9581
NLSA [64] 4$\times$ 44.99 32.59 0.9000 28.87 0.7891 27.78 0.7444 26.96 0.8109 31.27 0.9184
IPT [10] 4$\times$ 115.63 32.64 - 29.01 - 27.82 - 27.26 - - -
@@
Appendix E Experimental Results
E.1 Full Quantitative Results
Due to the limited space in the main manuscript, we only report a part of the experimental result. We show the quantitative experimental of grayscale image JPEG CAR in Tab. 11. The full results of image SR in Tab. 12, individually. These quantitative comparisons further validate the effectiveness of the proposed SemanIR.
E.2 More visual Results
@@
JPEG Compression Artifact Removal For JPEG compression artifact removal, the visual results for grayscale and color images on the Urban100 dataset are shown in Fig. 15 and Fig. 16, respectively. The proposed method achieves state-of-the-art performance in removing the blocking artifacts in the input images.
Figure 15: Visual comparison of grayscale JPEG CAR on Urban100 dataset. Best viewed by zooming.

### v2

@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
We thank Danda Pani Paudel for the valuable discussions on this work. This work was partially supported by Shenzhen Innovation in Science and Technology Foundation for The Excellent Youth Scholars (No. RCYX20231211090248064), the National Natural Science Foundation of China (No. 62203476), the MUR PNRR project FAIR (PE00000013) funded by the NextGenerationEU, the PRIN project CREATIVE (Prot. 2020ZSL9F9), the EU Horizon project ELIAS (No. 101120237), and the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure).
@@ A.3 Efficiency Analysis
$\mathcal{O}(6\times[4HWC^{2}+2kHWC]+(HW)^{2}C)$ (5)
@@
$\mathcal{O}(6\times[4HWC^{2}+2(M)^{2}HWC]-(6\times[4HWC^{2}+2kHWC]+(HW)^{2}C))$ (6)
$=\mathcal{O}((12M^{2}-12k-HW)HWC)$
In the last equation, to provide further clarity, let’s consider a common setting where the window size $M=7$ and the patch size is 16. The height $H$ and the width $W$ of the feature map are 64. The number of pixels within the window is approximately (7 $\times$ 7) $\times$ (16 $\times$ 16). In the proposed SemanIR, the k value is set to 512 or randomly sampled from [64, 128, 256, 384, 512]. To this end, we have:
$\mathcal{O}((12M^{2}-12k-HW)HWC)$ $=\mathcal{O}(12\times(7\times 7)\times(16\times 16)-12\times 512-64\times 64)$ (7)
$=\mathcal{O}(150528-6144-4096)>>0$
This shows that the complexity is significantly greater than zero.
@@ A.4 Training Details
Our method explores 6 IR tasks, and the training settings vary slightly for each task. These differences encompass the architecture of the proposed SemanIR, variations in the choice of the optimizer, and loss functions. Each experiments are conducted on 4 NVIDIA Tesla V100 32G GPUs.
@@ Appendix D More Ablation Analyses
The Impact of the $k$ in Key-Semantic Dictionary Construction under Various IR Tasks. To explore how the $k$ value of top-k will affect the IR performance of the proposed SemanIR. We conduct exhaustive experiments on JPEG compression artifact reduction for color images under different QF values (i.e., QF = [10, 20, 30, 40]), image denoising for both color and grayscale images under different noise levels (i.e., $\sigma$ = [15, 25, 50]), as well as image SR under different scales (i.e., 2$\times$, 3$\times$, 4$\times$) with the proposed SemanIR. Note that all the experiments for each IR task are conducted under two kinds of top-k settings, i.e., (i) $k$ was randomly sampled from the range [64, 512] during the overall training phase, and (ii) $k$ was held constant at 512 throughout the training phase. For inference, $k$ was configured to the specified value for both settings.
The results of the JPEG CAR in terms of the hyper-parameters $k$ under different training settings during inference for color image are shown in Fig. 8. It is clear that for the color JPEG CAR task when $k$ is set to 64 during inference, there is a huge performance cat between the random top-k setting and the fixed top-k setting. In addition, the fixed top-k setting performs well or sometimes even a bit better than the random top-k setting only when $k$ is also set to the same number (i.e., 512). With the decrease of $k$ during inference for the fixed top-k setting, the PSNR drops largely marginally for all the datasets under every kind of degraded QF factor.
Figure 12: Ablation study on the impact of $k$ for Image SR with SemanIR-B on BSD100 (a-c), Manga109 (d-f), Set5 (g-i), Set14 (i-l) and Urban100 (m-o) datasets with $scale$ = (2$\times$, 3$\times$, and 4$\times$).
@@
Appendix E More visual Results
@@
JPEG Compression Artifact Removal For JPEG compression artifact removal, the visual results for color images on the Urban100 dataset are shown in Fig. 15. The proposed method achieves state-of-the-art performance in removing the blocking artifacts in the input images.